ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 90 Zinchenko Olga, Karachai-Cherkes Technological Institute, Russia The extreme graphs problems The admissible decisions are presented in kind of fractal trees (wich appear by separate kind of fractal graphs) in such areas as the sequential analysis, cooperative decision making and etc. The report is devoted to problem of finding the dyadic fractal tree in the given graph. The known bounded degree tree is most relation to the our problem. This problem is NP-complete. The base of result of present scientific paper is the construction and substation of effective statically algorithm wich finds almost always a pre-fractal dyadic tree with polynomical computional compexity. Reference: Perepelitsa V. A., Pinchuk V. P., Sergeeva L. N. , Pozdnjikova A. J. Fractal Graphs and their Properties. ICM 97, 26. Februar bis l. März 1997, Weimar, banhous - Universität Weimar.